### v1

# Removing quasiconformal orbits

Abstract.
We show that generic continuous linear cocycles over shifts and other zero-dimensional systems admit no quasiconformal orbits, thus providing a partial answer to a question of Nassiri, Rajabzadeh, and Reshadat. The proof relies on a new result about towers for homeomorphisms of zero-dimensional spaces, which may be of independent interest.
2020 Mathematics Subject Classification:
37D30; 37B02, 37B20
1. Introduction
Linear cocycles model compositions of linear maps that are driven by an underlying dynamical system. They have been studied from various perspectives and are useful in diverse contexts: see the textbooks [Ar, BarP, CoK, DaF, J, Vi]. This paper deals with discrete-time continuous linear cocycles in the topological (continuous) setting. We now provide precise a precise definition.
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $X$. Let $\mathbb{F}$ be the either the field $\mathbb{R}$ or the field $\mathbb{C}$. Let $d$ be a positive integer. A $d$-dimensional continuous linear cocycle over $T$ is a continuous map $(x,n)\in X\times\mathbb{Z}\mapsto F^{(n)}(x)\in\mathrm{GL}(d,\mathbb{F})$ such that
(1) $F^{(m+n)}(x)=F^{(m)}(T^{n}x)F^{(n)}(x)\quad\text{\emph{(cocycle identity)}.}$) .
Note that the cocycle is uniquely determined by the map $F\coloneq F^{(1)}$. For this reason, we also refer to the pair $(T,F)$ as a linear cocycle. The set of linear cocycles over $T$ can be identified with the space $C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$, which we endow with the complete metric
(2) $d(F,G)\coloneqq\sup_{x\in X}\left(\|F(x)-G(x)\|+\|F(x)^{-1}-G(x)^{-1}\|\right)\,,$
For definiteness, $\|A\|$ will denote the Euclidean-induced operator norm of a matrix $A$; equivalently, $\|A\|=\sigma_{1}(A)$, where $\sigma_{1}(A)\geq\cdots\geq\sigma_{d}(A)$ denote the singular values of $A$.
Let $(T,F)$ be a continuous linear cocycle as above, and fix an integer $k$ with $0<k<d$. Let us say that the cocycle admits an exponential gap between $k$-th and $k+1$-th singular values if there exist positive constants $c$ and $\tau$ such that, for all $x\in X$ and $n>0$,
(3) $\sigma_{k}(F^{(n)}(x))>ce^{\tau n}\sigma_{k+1}(F^{(n)}(x))\,.$
As proved in [BoG], this condition is satisfied if and only if the cocycle admits a dominated splitting with a dominating bundle of dimension $k$ – we refer to the article for the exact definition. We note that dominated splittings play an important role in many problems in dynamics: see [BoG, BoPS, NRR, S] and references therein.
The condition number of a matrix $A\in\mathrm{GL}(d,\mathbb{F})$ is defined as
(4) $\kappa(A)\coloneqq\|A\|\,\|A^{-1}\|=\frac{\sigma_{1}(A)}{\sigma_{d}(A)}\,.$
Note that $\kappa(A)\geq 1$, with equality if and only if $A$ preserves the standard inner product up to a scalar factor. Following Nassiri, Rajabzadeh, and Reshadat [NRR], given a continuous linear cocycle as above, we say that a point $x\in X$ has a quasiconformal orbit if
(5) $\sup_{n\in\mathbb{Z}}\kappa(F^{(n)}(x))<\infty\,.$
It is clear that a cocycle admitting a dominated splitting cannot have quasiconformal orbits. It is also clear that the converse statement is false, that is, a cocycle may fail to have a dominated splitting and at the same time also fail to have quasiconformal orbits: consider for example a cocycle generated by non-trivial unipotent matrices. Nevertheless, in some situations, the converse statement is true up to perturbation: see [NRR, Theorem A] (see also [Bon, Proposition 2.1], [BoR, Theorem 3] for related results).
Given a continuous cocycle $(T,F)$ taking values in the group $\mathrm{GL}(2,\mathbb{R})$, if $p\in X$ is a periodic point of period $\ell$ and the matrix $F^{(\ell)}(p)$ has non-real eigenvalues, then the orbit of $p$ is quasiconformal. Furthermore, this property will persist under sufficiently small perturbations of $F$ in the $C^{0}$ topology. Thus, existence of quasiconformal orbits can be $C^{0}$-robust. However, this reasoning breaks down if the dimension is at least $3$, or if the field is $\mathbb{C}$.
If the base dynamics is a full shift, Nassiri, Rajabzadeh, and Reshadat [NRR, Theorem B] show that for any dimension $d\geq 2$, Hölder continuous $\mathrm{GL}(d,\mathbb{R})$-cocycles possessing quasiconformal orbits form a set of nonempty interior with respect to the appropriate Hölder topology. They also ask whether this robustness is possible in the $C^{0}$ topology (excluding the trivial $2$-dimensional case explained above): see [NRR, Question 8.2]. The goal of the present note is to answer this question in the negative, at least under some additional hypotheses on the topological dynamical system $(X,T)$.
Let us recall that a topological space has zero topological dimension if every open cover can be refined to an open partition. In this paper, we work with compact metric spaces, in which case having zero topological dimension is equivalent to the existence of a basis of the topology formed by clopen sets, among several other characterizations: see [Coo, Theorem 2.9.1] and [P, Theorem 2.10].
Our result is as follows:
Theorem 1.1.
Let $X$ be a compact metric space of zero topological dimension. Let $T\colon X\to X$ be a homeomorphism with finitely many periodic points of any given period. Let $\mathbb{F}$ be either $\mathbb{C}$ or $\mathbb{R}$, and let $d\geq 2$. Additionally, assume that $d\geq 3$ if $\mathbb{F}=\mathbb{R}$ and $T$ has periodic points. Let $\mathcal{QC}$ be the set of $F\in C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$ such that the corresponding cocycle admits quasiconformal orbits. Then the set $\mathcal{QC}$ is meager in $C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$.
Theorem 1.1 applies, for instance, to any subshift on a finite alphabet.
Remark 1.2.
The conclusions of Theorem 1.1 hold true if the set $X$ is finite, as an elementary linear algebra argument shows, or if the dynamics $(X,T)$ is minimal and the field $\mathbb{F}$ is $\mathbb{R}$, by [NRR, Corollary 3.6, Remark 3.7].
Remark 1.3.
Following [NRR], we say that a cocycle $(T,F)$ has a bounded orbit if $\sup_{n\in\mathbb{Z}}\|F^{(n)}(x)\|<\infty$ for some $x\in X$. It readily follows from Theorem 1.1 that, under the same hypotheses, the $\mathrm{SL}(d,\mathbb{F})$-valued cocycles admitting bounded orbits forms a meager subset of $C^{0}(X,\mathrm{SL}(d,\mathbb{F}))$. By contrast, the existence of bounded vectorial orbits (that is, the existence of $x\in X$ and $v\in\mathbb{R}^{d}\smallsetminus\{0\}$ such that $\sup_{n\in\mathbb{Z}}\|F^{(n)}(x)v\|<\infty$) is essentially equivalent to failure of uniform hyperbolicity, by results of Mañé and Sacker–Sell [ChL, Theorems 6.51, 6.52] (see also [DaF, Theorem 3.8.2] for the $\mathrm{SL}(2,\mathbb{R})$ case). Therefore, existence of bounded vectorial orbits may have non-empty interior in $C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$.
To motivate the proof of the theorem, note that breaking the quasiconformality of any individual orbit would be straightforward if we were allowed to perturb the matrices along that orbit independently. The difficulty lies in achieving this simultaneously for all orbits through a coherent and continuous perturbation of the cocycle. As one might anticipate, our approach relies on a tower construction, combined with a simple Baire argument.
Towers are used in dynamics in innumerable ways, starting with the Rokhlin–Halmos lemma (see e.g. [K]). Towers are particularly useful in low-regularity constructions like the one in the present paper.
Towers are inimical to periodic points: a tower of height $n$ cannot contain any periodic points of period $n$ or less. The Rokhlin–Halmos lemma assumes that periodic points form a set of measure zero, and thus can be ignored. In our setting, as we aim for a uniform and not an almost-everywhere statement, we cannot disregard the periodic points. Furthermore, the nature of the problem makes periodic orbits somewhat troublesome.
We overcome the difficulties above by proving a new type of tower lemma. Informally, we find partitions of the whole space $X$ by (clopen) towers with the following crucial property. Short towers, meaning those of height less than a given threshold $N$, actually behave as if they were tall: any orbit, upon first entering a short tower, is captured and must spend at least $N$ time units in the tower before exiting. This is the content of Theorem 4.1 below, which is a result in topological dynamics (without any direct relation to linear cocycles), and may be of independent interest.
The remainder of this paper is organized as follows: Section 2 discusses the basics of towers and castles. Section 3 establishes a capturing property around periodic points. The proof of the new tower lemma is given in Section 4. These three sections focus exclusively on topological dynamics. In Section 5 we return to linear cocycles and prove Theorem 1.1. In the final Section 6, we briefly discuss the possibility of improving our theorems and provide references for some related results.
[…]
2. Towers and castles
Definition 2.1.
Let $T\colon X\to X$ be a bijective map. Suppose $\ell\geq 1$ and $B\subseteq X$ are such that $B\cap T^{-i}B=\varnothing$ for all $i\in\ldbrack 0,\ell-1\rdbrack$. Then the set
(6) $K=\bigsqcup_{i=0}^{\ell-1}T^{i}B$
is called a tower with base $B$ and height $\ell$ with respect to the dynamics $T$, and each of the (disjoint) sets $B,TB,\dots,T^{\ell-1}B$ is called a floor of the tower. A castle is any disjoint union of towers, and its base is the union of their bases. When a castle covers the whole space we call it a castle partition.
Castles are also called skyscrapers or multitowers in the literature. We follow the terminology of [P, p. 39].
In this paper, we are interested in constructing castles composed of finitely many towers. For that reason, the following defintion will prove itself useful:
Definition 2.2.
Let $T\colon X\to X$ be a bijective map. A subset $E\subseteq X$ is called feedback set with respect to $T$ if there exists $n_{0}>0$ such that every segment of orbit of length $n_{0}$ hits $E$, that is,
(7) $\bigcup_{n=0}^{n_{0}-1}T^{-n}E=X\,.$
Equivalently, $E$ is a feedback set if and only if for every $x\in X$, the set of times $n\in\mathbb{Z}$ such that $T^{n}x\in E$ is syndetic (i.e., has bounded gaps). A feedback set is also called a marker set: see [Do, p. 98, Definition 2]. A related but distinct notion in measurable dynamics is that of a sweep-out set: see [AlP, p. 162]. Our terminology is motivated by the following example:
Example 2.3.
Suppose $T$ is a two-sided subshift of finite type (not necessarily of memory $1$). Then $T$ can be represented as the edge shift of a higher edge graph $G$: see [LM, Theorem 2.3.2]. Suppose $F$ is a set of edges of $G$ which is a feedback set in the sense of graph theory [BanG, Chapter 15], that is, the removal of those edges makes the graph acyclic. Let $E\subseteq X$ be the union of all cylinders corresponding to the edges in $F$; then $E$ is a feedback set in the sense of Definition 2.2. The same construction applies to any homeomorphism of a zero-dimensional space, using graph covers [Bor].
The following is a finitary version of the Kakutani–Rokhlin skyscraper construction:
Lemma 2.4.
Let $T\colon X\to X$ be a bijective map and let $E\subseteq X$ be a feedback set. Then there exists a partition of $X$ into finitely many towers forming a castle with base $E$. Additionally, if $T$ is a homeomorphism of a topological space, and $E$ is a clopen subset, then the castle partition above can be chosen in such a way that each floor of each tower is a clopen set.
Proof.
Take $n_{0}>0$ such that property (7) holds. For each $\ell\in\ldbrack 1,n_{0}\rdbrack$, let
(8) $B_{\ell}\coloneqq E\cap T^{-\ell}(E)\smallsetminus\bigcup_{i=1}^{\ell-1}T^{-i}(E)\,,$
that is, the set of points in $E$ whose first return to $E$ occurs at time $\ell$. These sets form a partition of $E$. Furthermore, each set $B_{\ell}$ is the base of a tower of height $\ell$, and those towers form a partition of $X$. If $T$ is a homeomorphism and $E$ is a clopen set, then each set $B_{\ell}$ is clopen. ∎
The following is a simple but useful observation:
Lemma 2.5.
Suppose $T\colon X\to X$ is a homeomorphism of a compact metric space $X$. An open set $E\subseteq X$ is a feedback set if and only if every (two-sided) orbit hits $E$, that is,
(9) $\bigcup_{n\in\mathbb{Z}}T^{-n}E=X\,.$
Proof.
Immediate. ∎
Suppose we want to find a castle partition of the space $X$ into uniformly high towers. There is an obstruction: $T$ cannot have periodic points of small period. The following Lemma shows that this obvious necessary condition is also sufficient, at least when the ambient space is zero-dimensional:
Lemma 2.6 (Downarowicz).
Let $X$ be a compact metric space of zero topological dimension and let $T\colon X\to X$ be a homeomorphism. Let $N>0$ and suppose that $T$ has no periodic points of period less than $N$. Then there exists a castle partition of $X$ into finitely many towers of height $\geq N$ whose floors are clopen sets.
Downarowicz original statement [Do, p. 98, Lemma 1] yields the conclusion of Lemma 2.6 under the stronger hypothesis that $T$ has no periodic points at all; nevertheless his argument proves the statement above. Downarowicz also allows $T$ to be non-invertible, which requires extra care. For the convenience of the reader we will provide a proof of the statement we need.
Proof of Lemma 2.6.
Suppose $T\colon X\to X$ is a homeomorphism without periodic points of period less than $N$, where $X$ is a zero-dimensional compact metric space $X$. Then each point in $X$ admits a clopen neighborhood $V$ which is disjoint from $TV,\dots,T^{N-1}V$. By compactness, we can cover $X$ by finitely many neighborhoods $V_{1},\dots,V_{m}$ of this type. Define a nested sequence of sets $B_{0}\subseteq B_{1}\subseteq\cdots\subseteq B_{m}$ recursively as follows: $B_{0}\coloneqq\varnothing$ and
(10) $B_{j+1}$ $\coloneqq B_{j}\cup(V_{j+1}\smallsetminus E_{j})\,,\quad\text{where}$
(11) $E_{j}$ $\coloneqq\bigcup_{n=-N+1}^{N-1}T^{n}B_{j}\,.$
We claim that
(12) $B_{j}\cap T^{n}B_{j}=\varnothing\text{ for all $j\in\ldbrack 0,m\rdbrack$ and $n\in\ldbrack 0,N-1\rdbrack$.}$
The claim is proved by induction on $j$: it is clearly true for $j=0$, and the induction step goes as follows:
(13) $B_{j+1}\cap T^{n}B_{j+1}$ $=\big{(}B_{j}\cup(V_{j+1}\smallsetminus E_{j})\big{)}\cap T^{n}\big{(}B_{j}\cup(V_{j+1}\smallsetminus E_{j})\big{)}$
$\subseteq(B_{j}\cap T^{n}B_{j})\cup(B_{j}\smallsetminus T^{n}E_{j})\cup(E_{j}\smallsetminus T^{n}B_{j})\cup(B_{j}\cap T^{n}B_{j})$
[…]
This proves (12).
Next, note that for each $j\in\ldbrack 1,m\rdbrack$,
(14) $V_{j}\subseteq\underbrace{(V_{j}\smallsetminus E_{j-1})}_{\subseteq B_{j}}\cup\underbrace{E_{j-1}}_{\subseteq E_{j}}\subseteq E_{j}\,.$
It follows that
(15) $E_{m}=\bigcup_{j=1}^{m}E_{j}\supseteq\bigcup_{j=1}^{m}V_{j}=X\,.$
In particular, $B_{m}$ is a feedback set. Consider the Kakutani–Rokhlin castle partition of $X$ with base $B_{m}$, given by Lemma 2.4. Each tower has height at least $N$, by (12). Since $B_{m}$ is a clopen set, all floors of all towers are clopen sets. ∎
3. Capturing property and periodic orbits
We now introduce the following crucial notion:
Definition 3.1.
Let $T$ be a homeomorphism of a compact metric space $X$, and let $N$ be a positive integer. A set $E\subseteq X$ is called $N$-capturing with respect to $T$ if
(16) $E\smallsetminus TE\subseteq\bigcap_{n=0}^{N-1}T^{-n}E\,.$
In other words, the set $E$ is $N$-capturing if every orbit, upon first entering $E$, remains in $E$ at least $N-1$ additional time units before exiting. Another formulation is follows: $E$ is $N$-capturing if and only if
(17) $\forall x\in E\ \exists m,n\geq 0\text{ s.t.\ }m+n+1\geq N\text{ and }\forall i\in\ldbrack-m,n\rdbrack,\ T^{i}x\in E\,.$
In particular, $E$ is also $N$-capturing with respect to $T^{-1}$.
Every tower of height $\ell\geq N$ is an $N$-capturing set. (On the other hand, a tower of height $\ell<N$ which is a $N$-capturing set is actually $\ell\lceil N/\ell\rceil$-capturing). Thus, non-periodic points have $N$-capturing neighborhoods, where $N$ can be chosen arbitrarily large. It turns out that the same is true for periodic points, as a consequence of the following lemma:
Lemma 3.2.
Let $X$ be a compact metric space and let $T\colon X\to X$ be a homeomorphism. Suppose $p$ is a periodic point of least period $\ell$. Let $N$ be a positive integer and let $U$ be a neighborhood of $p$. Then there exists a neighborhood $B$ of $p$ with $B\subseteq U$ such that the sets $B,TB,\dots,T^{\ell-1}B$ are disjoint, their union
(18) $K\coloneqq\bigsqcup_{i=0}^{\ell-1}T^{i}B\,,$
is a $N$-capturing set, and its complement $K^{\mathsf{c}}$ is also a $N$-capturing set. Furthermore, if $X$ is zero-dimensional, then $B$ can be chosen as a clopen set.
The statement above should be evident if $p$ is a hyperbolic periodic point of a diffeomorphism, for instance: see Fig. 1.
We will prove Lemma 3.2 initially for fixed points ($\ell=1$), and later deduce the general statement from this particular case.
Figure 1. An $N$-capturing neighborhood $B$ of a hyperbolic fixed point. If $B$ is taken small enough, then the complement $B^{\mathsf{c}}$ is also $N$-capturing.
Proof of Lemma 3.2 when $\ell=1$.
Suppose $p$ is a fixed point of $T$. Given $N>0$ and a neighborhood $U$ of $p$, take a smaller neighborhood $V$ such that
(19) $\bigcup_{i=-2N+1}^{2N-1}T^{-i}V\subseteq U\,.$
Define sets
(20) $V$ $\coloneqq\bigcap_{i=0}^{N-1}T^{-i}V\,,\qquad V\coloneqq\bigcap_{i=0}^{N-1}T^{i}V=T^{N}V\,,\quad\text{and}$
(21) $C$ $\coloneqq\bigcup_{j=0}^{N-1}T^{j}V=\bigcup_{j=0}^{N-1}T^{-j}V\,.$
We have
(22) $C\smallsetminus TC=\bigcup_{j=0}^{N-1}T^{j}V\smallsetminus\bigcup_{j=1}^{N}T^{j}V\subseteq V=\bigcap_{j=0}^{N-1}T^{-j}(T^{j}V)\subseteq\bigcap_{j=0}^{N-1}T^{-j}(C)\,,$
that is, the set $C$ is $N$-capturing.
Define other sets
(23) $B$ $\coloneqq C\cup D\,,\quad\text{where}$
(24) $D$ $\coloneqq\bigcup_{\begin{subarray}{c}i,j\geq 1\\ i+j\leq N\end{subarray}}(T^{-i}V\cap T^{j}V)\,.$
Note that
(25) $T^{-1}D=\bigcup_{\begin{subarray}{c}i\geq 2,\,j\geq 0\\ i+j\leq N\end{subarray}}(T^{-i}V\cap T^{j}V)\subseteq V\cup D\subseteq B\,.$
Then,
(26) $B\smallsetminus TB$ $\subseteq C\smallsetminus TB$ (since $B=C\cup D$ and $D\subseteq TB$)
(27) $\subseteq C\smallsetminus TC$ (since $B\supseteq C$)
(28) $\subseteq\bigcap_{n=0}^{N-1}T^{-n}C$ (since $C$ is $N$-capturing)
(29) $\subseteq\bigcap_{n=0}^{N-1}T^{-n}B$ (since $B\supseteq C$),
proving that $B$ is $N$-capturing.
We will check that $B^{\mathsf{c}}$ is also $N$-capturing. A reasoning analogous to (25) shows that $TD\subseteq B$. On the other hand,
(30) $C\smallsetminus T^{-1}C=\bigcup_{j=0}^{N-1}T^{-j}V\smallsetminus\bigcup_{j=1}^{N}T^{-j}V\subseteq V\,.$
Then the following chain of inclusions holds:
(31) $B^{\mathsf{c}}\smallsetminus TB^{\mathsf{c}}$ $=TB\smallsetminus B$
(32) $\subseteq TC\smallsetminus B$ (since $B=C\cup D$ and $TD\subseteq B$)
(33) $=T(C\smallsetminus T^{-1}C)\smallsetminus D$
(34) $\subseteq TV\smallsetminus D$ (by (30)).
It follows from (23) that
(35) $D\supseteq TV\cap\bigcup_{i=1}^{N-1}T^{-i}V\,.$
Resuming from (34), we have the following inclusions:
(36) $B^{\mathsf{c}}\smallsetminus TB^{\mathsf{c}}$ $\subseteq TV\smallsetminus D$
(37) $\subseteq\bigcap_{i=1}^{N-1}T^{-i}V^{\mathsf{c}}$ (by (35))
(38) $\subseteq\bigcap_{i=1}^{N-1}T^{-i}B^{\mathsf{c}}$ (since $V\subseteq C\subseteq B$),
proving that $B^{\mathsf{c}}$ is $N$-capturing.
Condition (19) ensures that $B\subseteq U$. Finally, if $X$ is zero-dimensional, then $V$ can be chosen as a clopen set, and it follows that $B$ is clopen. ∎
Proof of Lemma 3.2 for arbitrary $\ell$.
Suppose $p$ is a periodic point of least period $\ell$. Reducing the neighborhood $U$ if necessary, we can assume that $U$ is disjoint from $TU,\dots,T^{\ell-1}U$. Given $N>0$, let $m\coloneqq\lceil N/\ell\rceil$. Consider the following smaller neighborhood of $p$:
(39) $V\coloneqq\bigcap_{j=0}^{m}T^{-\ell j}U\,.$
Applying the previously proved version of the Lemma (with $T^{\ell}$, $V$, and $m$ in the respective places of $T$, $U$, and $N$), we obtain a neighborhood $B$ of $p$ such that $B\subseteq V$, both $B$ and $B^{\mathsf{c}}$ are $m$-capturing with respect to $T^{\ell}$, and $B$ is clopen if $\dim X=0$. Let $K\coloneqq\bigsqcup_{i=0}^{\ell-1}T^{i}B$. Note that
(40) $B\subseteq\bigcap_{i=0}^{\ell-1}T^{-i}K\,.$
We will check that both sets $K$ and $K^{\mathsf{c}}$ are $\ell m$-capturing (and a fortiori $N$-capturing) with respect to $T$. We have
(41) $K\smallsetminus TK=\bigcup_{i=0}^{\ell-1}T^{i}B\smallsetminus\bigcup_{i=1}^{\ell}T^{i}B\\ \subseteq B\smallsetminus T^{\ell}B\subseteq\bigcap_{j=0}^{m-1}T^{-\ell j}B\,,$
since $B$ is $m$-capturing with respect $T^{\ell}$. Using (40), we obtain
(42) $K\smallsetminus TK\subseteq\bigcap_{j=0}^{m-1}\bigcap_{i=0}^{\ell-1}T^{-\ell j-i}K=\bigcap_{n=0}^{\ell m-1}T^{-n}K\,,$
showing that $K$ is $\ell m$-capturing.
Assume for a contradiction that $K^{\mathsf{c}}$ is not $\ell m$-capturing, that is, there exists $x\in K^{\mathsf{c}}\smallsetminus TK^{\mathsf{c}}$ such that $T^{n}x\in K$ for some $n\in\ldbrack 0,\ell m-1\rdbrack$. Then
(43) $x\in TK\smallsetminus K=\bigcup_{i=1}^{\ell}T^{i}B\smallsetminus\bigcup_{i=0}^{\ell-1}T^{i}B\\ \subseteq T^{\ell}B\smallsetminus B\,.$
Use Euclidean division and write $n=\ell k+r$ with $k\in\ldbrack 0,m-1\rdbrack$ and $r\in\ldbrack 0,\ell-1\rdbrack$. Then,
(44) $T^{n}x\in T^{n+\ell}B\subseteq T^{n+\ell}V=T^{\ell(k+1)+r}V\subseteq T^{r}U\,,$
by (39). The point $T^{n}x$ belongs to $K$, that is, it belongs to one of the sets $B,\dots,T^{\ell-1}B$, which are respectively contained in the disjoint sets $U,\dots,T^{\ell-1}U$. Hence the point $T^{n}x$ must belong to $T^{r}B$. That is, $T^{\ell k}x\in B$. However, $x\in B^{\mathsf{c}}$, therefore contradicting the fact that $B^{\mathsf{c}}$ is $m$-capturing with respect to $T^{\ell}$. ∎
4. Castles with capturing towers
Our new tower lemma announced in the introduction is stated as follows (where the capturing property is as in Definition 3.1):
Theorem 4.1.
Let $X$ be a compact metric space of zero topological dimension and let $T\colon X\to X$ be a homeomorphism. Let $N>0$ and suppose that the set of periodic points of period less than $N$ is finite. Let $\delta>0$. Then there exists a castle partition of $X$ into finitely many $N$-capturing towers whose floors are clopen sets of diameter less than $\delta$.
Proof.
If $T$ has no periodic points of period less than $N$, then by Lemma 2.6 there exists a partition of $X$ into finitely many towers of height at least $N$ whose floors are clopen sets. These towers are automatically $N$-capturing. Furthermore, we can slice the towers so that each floor has diameter less than $\delta$. This proves Theorem 4.1 in this particular case.
Next, assume that $T$ has some periodic orbits of period less than $N$, but only finitely many of them: say $\mathcal{O}_{1},\dots,\mathcal{O}_{s}$ (all distinct), with corresponding least periods $\ell_{1},\dots,\ell_{s}$. Let
(45) $\delta_{0}\coloneqq\min_{i\neq j}d(\mathcal{O}_{i},\mathcal{O}_{j})\,.$
(If $s=1$, then $\delta_{0}\coloneqq\infty$.) Reducing $\delta$ if necessary, we can assume that $\delta<\delta_{0}/2$ and
(46) $\forall x,y\in X,\text{ if }d(x,y)<\delta,\text{ then }\min_{n\in\ldbrack 1,N\rdbrack}d(T^{n}x,T^{n}y)<\frac{\delta_{0}}{2}\,.$
Applying Lemma 3.2, for each $i$ we can find a clopen set $B_{i}$ such that the tower
(47) $K_{i}\coloneqq B_{i}\sqcup TB_{i}\sqcup\cdots\sqcup T^{\ell_{i}-1}B_{i}$
contains $\mathcal{O}_{i}$, is $N$-capturing, and its complement $K_{i}^{\mathsf{c}}$ is also $N$-capturing. Furthermore, $\operatorname{diam}B_{i}$ can be chosen sufficiently small so that each floor of $K_{i}$ has diameter less than $\delta$. In particular, the set $K_{i}$ is contained in the $\delta_{0}/2$-neighborhood of the orbit $\mathcal{O}_{i}$. It follows from (45) that the sets $K_{1},\dots,K_{s}$ are pairwise disjoint. Let
(48) $K\coloneqq\bigsqcup_{i=1}^{s}K_{i}\quad\text{and}\quad B\coloneqq\bigsqcup_{i=1}^{s}B_{i}\,.$
So, $K$ is a castle with base $B$. The set $K$ is $N$-capturing, since each of the towers $K_{i}$ is $N$-capturing.
We claim that the complement set $K^{\mathsf{c}}$ is also $N$-capturing. Indeed, assume this assertion is false, that is, there exists a point $x\in K$ such that $Tx\not\in K$ and $T^{n}x\in K$ for some $n\in\ldbrack 2,N\rdbrack$. Let $i$ and $j$ be such that $x\in K_{i}$ and $T^{n}x\in K_{j}$. Since $K_{i}^{\mathsf{c}}$ is $N$-capturing, we must have $j\neq i$. We have $d(x,\mathcal{O}_{i})<\delta$ and so (46) ensures that $d(T^{n}x,\mathcal{O}_{i})<\delta_{0}/2$. Furthermore, $d(T^{n}x,\mathcal{O}_{j})<\delta<\delta_{0}/2$, so the triangle inequality yields $d(\mathcal{O}_{i},\mathcal{O}_{j})<\delta_{0}$, which contradicting (45). This proves our claim: the set $K^{\mathsf{c}}$ is $N$-capturing.
Let $M$ be the maximal invariant set of $K^{\mathsf{c}}$, i.e.,
(49) $M\coloneqq\bigcap_{n\in\mathbb{Z}}T^{-n}(K^{\mathsf{c}})\,.$
Then $M$ is a (possibly empty) closed set and the restricted map $T|_{M}$ has no periodic points of period less than $N$. By Lemma 2.6, $M$ admits a finite castle partition
(50) $M=\bigsqcup_{j=1}^{t}\bigsqcup_{n=0}^{m_{j}-1}T^{n}(D_{j})\,,$
into towers of heights $m_{j}\geq N$ whose bases $D_{j}$ are clopen sets in the relative topology of $M$.
Next, we enlarge the castle (50): for each $i\in\ldbrack 1,t\rdbrack$, let $C_{j}$ be a clopen neighborhood of the compact set $D_{j}$. Choosing those neighborhoods small enough, the following union is disjoint:
(51) $\bigsqcup_{j=1}^{t}\bigsqcup_{n=0}^{m_{j}-1}T^{n}(C_{j})\eqcolon L\,;$
furthermore, the resulting castle $L$ is disjoint from $K$. Let $C\coloneqq\bigsqcup_{j=1}^{t}C_{j}$ be the base of the castle $L$ (see Fig. 2).
Figure 2. The two castles $K$ and $L$. The arrows denote possible excursions outside of $K\sqcup L$. Since the set $K^{\mathsf{c}}$ is $N$-capturing, excursions following the leftmost arrow take time at least $N$.
We claim that $B\sqcup C$ is a feedback set. By Lemma 2.5 it is sufficient to check the (two-sided) orbit of every point $x\in X$ hits $B\sqcup C$. If $x\not\in M$, then, by definition of the set $M$, the orbit of $x$ hits $K$ and hence hits $B$. On the other hand, if $x\in M$, then the orbit of $x$ never hits the invariant castle $M$, and so it hits its base, which is contained in $C$. This proves our claim; actually, this argument shows that $B\sqcup(C\cap T(L))$ is also a feedback set.
Define another clopen set:
(52) $E\coloneqq B\sqcup(T(L)\cap C)\sqcup(T(K)\smallsetminus K)\,.$
Then $E$ is a feedback set, since it contains the feedback set $B\sqcup(C\cap T(L))$. Let $Q$ be the Kakutani–Rokhlin castle with base $E$ given by Lemma 2.4; this is a partition of the whole space $X$ into finitely many towers with clopen floors. Note that each tower $K_{i}$ of the castle $K$ is contained in a tower of the same height $\ell_{i}$ of the new castle $Q$; indeed, if $x$ is in the base $B_{i}$ of the tower $K_{i}$, then $T^{\ell_{i}}(x)$ belongs to $B\sqcup(T(K)\smallsetminus K)$ and hence to $E$, the base of $Q$.
Therefore we can refine the castle $Q$ (by slicing the towers that intersect $K$) and obtain a castle $Q^{\prime}$ with the same base $E$ such that every tower of $K$ is also a tower of $Q^{\prime}$. Recall that each of these towers have heights less than $N$, are $N$-capturing, and have floors of diameters less than $\delta$. By further slicing of towers, we can assume that all floors of all towers of $Q^{\prime}$ have diameter less than $\delta$.
To complete the proof of the Theorem, it remains to check that each tower of $Q^{\prime}$ which is not a tower of $K$ has height at least $N$. Equivalently, for each point $x$ in the set $E\smallsetminus B=(T(L)\cap C)\sqcup(T(K)\smallsetminus K)$, the segment of orbit $Tx,T^{2}x,\dots,T^{N-1}x$ does not hit $E$.
First, consider the case where $x\in T(L)\cap C$. Then the segment of orbit above is contained in $L\smallsetminus C$ and thus in $E^{\mathsf{c}}$ (note that $(T(K)\smallsetminus K)\cap L\subseteq L\smallsetminus T(L)\subseteq C$).
Next, consider the remaining case where $x\in T(K)\smallsetminus K$. Consider the return time to $E$, that is, the least positive $n>0$ such that $T^{n}x\in E$. The point $T^{n}x$ cannot belong to $T(K)\smallsetminus K$, otherwise there would be an earlier hit to $B\subseteq E$. Therefore, $T^{n}x\in B\sqcup(T(L)\cap C)$. If $T^{n}x$ is in $B$, then the fact that the set $K^{\mathsf{c}}$ is $N$-capturing tells us that $n\geq N$, as desired. If, on the other hand, $T^{n}x$ is in $T(L)\cap C$, then the point $T^{n-1}x$ is in the top floor of a tower of $L$, and since $x\not\in L$, the segment of orbit $Tx,T^{2}x,\dots,T^{n-1}x$ must include a full traverse of that tower. In particular, $n$ is bigger than the height of the tower, which is at least $N$.
This completes the proof of the Theorem. ∎
5. Eradicating quasiconformal orbits
In this section, we use the previous results about castles to prove our main goal, Theorem 1.1. We need a few preliminaries in linear algebra.
The spectral radius and the eigenvalue condition number of a square matrix $A$ are defined respectively as:
(53) $\rho(A)$ $\coloneqq\max\{|\lambda|\;\mathord{:}\;\lambda\in\mathbb{C}\text{is an eigenvalue of $A$}\}\,,$
(54) $\kappa_{\mathrm{e}}(A)$ $\coloneqq\rho(A)\rho(A^{-1})\,.$
Note that $\rho(A^{n})=\rho(A)^{n}$ for every $n>0$ and $\kappa_{\mathrm{e}}(A^{n})=\kappa_{\mathrm{e}}(A)^{n}$ for every $n\neq 0$.
Lemma 5.1.
Let $\mathbb{F}$ be either $\mathbb{C}$ or $\mathbb{R}$ and let $d\geq 2$. Assume that $d\geq 3$ if $\mathbb{F}=\mathbb{R}$. Let $A_{0},\dots,A_{\ell-1}$ be a finite list of matrices in $\mathrm{GL}(d,\mathbb{F})$. Then, for every $\varepsilon>0$, there exists $\tilde{A}_{0},\dots,\tilde{A}_{\ell-1}\in\mathrm{GL}(d,\mathbb{F})$ such that $\|\tilde{A}_{j}-A_{j}\|\leq\varepsilon\|A_{j}\|$ for each $j\in\ldbrack 0,\ell-1\rdbrack$ and
(55) $\kappa_{\mathrm{e}}\big{(}\tilde{A}_{\ell-1}\cdots\tilde{A}_{0}\big{)}\geq(1+\varepsilon)^{\ell}\,.$
Proof.
Let $\lambda_{1},\dots,\lambda_{d}$ be the eigenvalues of the matrix $A_{\ell-1}\cdots A_{0}$, ordered so that $|\lambda_{1}|\geq\cdots\geq|\lambda_{d}|$ and furthermore $\lambda_{d-1}=\overline{\lambda_{d}}$ in the case that $\mathbb{F}=\mathbb{R}$ and $\lambda_{d}\in\mathbb{C}\smallsetminus\mathbb{R}$. If $\mathbb{F}=\mathbb{R}$ and $\lambda_{d}\in\mathbb{C}\smallsetminus\mathbb{R}$, let $\nu\coloneq 2$, otherwise let $\nu\coloneqq 1$. Let $V_{0}\subset\mathbb{F}^{d}$ be an $\nu$-dimensional eigenspace of $A_{\ell-1}\cdots A_{0}$ corresponding to the eigenvalue $\lambda_{d}$ if $\nu=1$, or to the eigenvalue pair $\lambda_{d-1},\lambda_{d}$ if $\nu=2$. For each $j\in\ldbrack 1,\ell-1\rdbrack$, define the $\nu$-dimensional subspace $V_{i}\coloneqq A_{i-1}\cdots A_{0}(V_{0})$. Let $\{e_{1},\dots,e_{d}\}$ denote the canonical basis of $\mathbb{F}^{d}$. For each $j\in\ldbrack 1,\ell-1\rdbrack$, take an orthogonal (if $\mathbb{F}=\mathbb{R}$) or unitary (if $\mathbb{F}=\mathbb{C}$) matrix $R_{j}$ such that $R_{j}V_{j}$ equals $\mathrm{span}(e_{d})$ if $\nu=1$ and $\mathrm{span}(e_{d-1},e_{d})$ if $\nu=2$. Let $R_{d}\coloneqq R_{0}$. For each $j\in\ldbrack 0,\ell-1\rdbrack$, let $\Delta_{j}\coloneqq R_{j+1}A_{j}R_{j}^{-1}$. Then $\Delta_{j}$ is a block triangular matrix:
(56) $\Delta_{j}\eqcolon\begin{pmatrix}B_{j}&0\\ D_{j}&C_{j}\end{pmatrix}\,,\quad\text{where $C_{j}$ is a $\nu\times\nu$ matrix.}$.
Multiplying these matrices, we get another block diagonal matrix
(57) $\Delta_{\ell-1}\cdots\Delta_{0}=\begin{pmatrix}B_{\ell-1}\cdots B_{0}&0\\ *&C_{\ell-1}\cdots C_{0}\end{pmatrix}$
Since $\Delta_{\ell-1}\cdots\Delta_{0}=R_{0}^{-1}A_{\ell-1}\cdots A_{0}R_{0}$, the submatrix $B_{\ell-1}\cdots B_{0}$ has eigenvalues $\lambda_{1},\dots,\lambda_{d-\nu}$, while the submatrix $C_{\ell-1}\cdots C_{0}$ has eigenvalues $\lambda_{d}$ if $\nu=1$ or $\lambda_{d-1},\lambda_{d}$ if $\nu=2$.
For each $j\in\ldbrack 0,\ell-1\rdbrack$, define matrices
(58) $\tilde{\Delta}_{j}\coloneqq\begin{pmatrix}(1+\varepsilon)B_{j}&0\\ D_{j}&C_{j}\end{pmatrix}\quad\text{and}\quad\tilde{A}_{j}\coloneq R_{j+1}^{-1}\tilde{\Delta}_{j}R_{j}\,.$
Note that
(59) $\|\tilde{A}_{j}-A_{j}\|=\|\tilde{\Delta}_{j}-\Delta_{j}\|=\left\|\left(\begin{smallmatrix}\varepsilon I_{d-\nu}&0\\ 0&0\end{smallmatrix}\right)\Delta_{j}\right\|\leq\varepsilon\|\Delta_{j}\|=\varepsilon\|A_{j}\|\,.$
Multiplying the block-triangular matrices $\tilde{\Delta}_{j}$, we obtain
(60) $\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0}=\begin{pmatrix}(1+\varepsilon)^{\ell}B_{\ell-1}\cdots B_{0}&0\\ *&C_{\ell-1}\cdots C_{0}\end{pmatrix}\,.$
This matrix has eigenvalues
(61) $(1+\varepsilon)^{\ell}\lambda_{1},\ \dots,\ (1+\varepsilon)^{\ell}\lambda_{d-1},\ \lambda_{d}$ if $\nu=1$, or
(62) $(1+\varepsilon)^{\ell}\lambda_{1},\ \dots,\ (1+\varepsilon)^{\ell}\lambda_{d-2},\ \lambda_{d-1},\ \lambda_{d}$ if $\nu=2$.
In any case,
(63) $\kappa_{\mathrm{e}}\big{(}\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0}\big{)}=\frac{(1+\varepsilon)^{\ell}|\lambda_{1}|}{|\lambda_{d}|}\geq(1+\varepsilon)^{\ell}\,.$
Since the matrix $\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0}$ is similar (i.e. conjugate) to $\tilde{A}_{\ell-1}\cdots\tilde{A}_{0}$, inequality (55) follows. ∎
[…]
(64) $\|A\|\geq\rho(A)\,,\quad\kappa(A)\geq\kappa_{\mathrm{e}}(A)\,.$
Since the norm is sub-multiplicative, so is the condition number:
(65) $\kappa(AB)\leq\kappa(A)\kappa(B)\,.$
In terms of the singular values $\sigma_{1}(A)\geq\cdots\geq\sigma_{d}(A)$ of the matrix $A$, we have $\|A\|=\sigma_{1}(A)$ and $\kappa(A)=\sigma_{1}(A)/\sigma_{d}(A)$.
Lemma 5.2.
Let $\mathbb{F}$ be either $\mathbb{C}$ or $\mathbb{R}$, let $d\geq 2$, and let $A_{0},\dots,A_{\ell-1}$ be a finite list of matrices in $\mathrm{GL}(d,\mathbb{F})$. Then, for every $\varepsilon>0$, there exists $\tilde{A}_{0},\dots,\tilde{A}_{\ell-1}\in\mathrm{GL}(d,\mathbb{F})$ such that $\|\tilde{A}_{j}-A_{j}\|\leq\varepsilon\|A_{j}\|$ for each $j\in\ldbrack 0,\ell-1\rdbrack$ and
(66) $\kappa\big{(}\tilde{A}_{\ell-1}\cdots\tilde{A}_{0}\big{)}\geq(1+\varepsilon)^{\ell}\,.$
Proof.
Let $v_{0}\in\mathbb{F}^{d}$ be a unit singular vector of the matrix $A_{\ell-1}\cdots A_{0}$ corresponding to the least singular value, that is,
(67) $\|v_{0}\|=1\quad\text{and}\quad\|A_{\ell-1}\cdots A_{0}v_{0}\|=\sigma_{d}(A_{\ell-1}\cdots A_{0})\,.$
For each $j\in\ldbrack 1,\ell\rdbrack$, define the unit vector
(68) $v_{j}\coloneqq\frac{A_{i-1}\cdots A_{0}v_{0}}{\|A_{i-1}\cdots A_{0}v_{0}\|}$
and take an orthogonal or unitary matrix $R_{j}$ such that $R_{j}v_{j}=e_{d}$, where $e_{d}\coloneqq(0,\dots,0,1)\in\mathbb{F}^{d}$. For each $j\in\ldbrack 0,\ell-1\rdbrack$, let $\Delta_{j}\coloneqq R_{j+1}A_{j}R_{j}^{-1}$. Then $\Delta_{j}$ is a block triangular matrix:
(69) $\Delta_{j}\eqcolon\begin{pmatrix}B_{j}&0\\ D_{j}&c_{j}\end{pmatrix}\,,\quad\text{where $B_{j}$ is a $(d-1)\times(d-1)$ matrix.}$.
By the SVD theorem, the product $\Delta_{\ell-1}\cdots\Delta_{0}$ has the same singular values as $A_{\ell-1}\cdots A_{0}$ and is a block diagonal matrix
(70) $\Delta_{\ell-1}\cdots\Delta_{0}=\begin{pmatrix}B_{\ell-1}\cdots B_{0}&0\\ 0&c_{\ell-1}\cdots c_{0}\end{pmatrix}$
with
(71) $\sigma_{d-1}(B_{\ell-1}\cdots B_{0})=\sigma_{d-1}(A_{\ell-1}\cdots A_{0})\geq\sigma_{d}(A_{\ell-1}\cdots A_{0})=|c_{\ell-1}\cdots c_{0}|\,.$
Analogously to (58), for each $j\in\ldbrack 0,\ell-1\rdbrack$ we define matrices
(72) $\tilde{\Delta}_{j}\coloneqq\begin{pmatrix}(1+\varepsilon)B_{j}&0\\ D_{j}&c_{j}\end{pmatrix}\quad\text{and}\quad\tilde{A}_{j}\coloneqq R_{j+1}^{-1}\tilde{\Delta}_{j}R_{j}\,.$
By the previous computation (59), we have $\|\tilde{A}_{j}-A_{j}\|\leq\varepsilon\|A_{j}\|$. Multiplying the block-triangular matrices $\tilde{\Delta}_{j}$, we obtain
(73) $\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0}=\begin{pmatrix}(1+\varepsilon)^{\ell}B_{\ell-1}\cdots B_{0}&0\\ *&c_{\ell-1}\cdots c_{0}\end{pmatrix}\,.$
This matrix expands the vector $e_{d}$ by a factor $|c_{\ell-1}\cdots c_{0}|$; in particular,
(74) $\sigma_{d}(\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0})\leq|c_{\ell-1}\cdots c_{0}|\,.$
On the other, the same matrix expands every vector orthogonal to $e_{d}$ by a factor at least $(1+\varepsilon)^{\ell}\sigma_{d-1}(B_{\ell-1}\cdots B_{0})$. It follows from the Courant-Fisher principle [HoJ, p. 451, eq. (7.3.10)] that
(75) $\sigma_{d-1}(\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0})\geq(1+\varepsilon)^{\ell}\sigma_{d-1}(B_{\ell-1}\cdots B_{0})\,.$
Therefore,
(76) $\kappa(\tilde{A}_{\ell-1}\cdots\tilde{A}_{0})=\kappa(\tilde{\Delta}_{\ell-1}\cdots\tilde{\Delta}_{0})\geq\frac{(1+\varepsilon)^{\ell}\sigma_{d-1}(B_{\ell-1}\cdots B_{0})}{|c_{\ell-1}\cdots c_{0}|}\geq(1+\varepsilon)^{\ell}\,.\qed$
We are ready to prove our main result.
Proof of Theorem 1.1.
Assume $X$ is a zero-dimensional compact metric space and that $T\colon X\to X$ is a homeomorphism with finitely many periodic points of any given period. Let $\mathbb{F}=\mathbb{C}$ or $\mathbb{R}$. If $\mathbb{F}=\mathbb{R}$ and $T$ has periodic points, assume that $d\geq 3$; otherwise, assume that $d\geq 2$. For each $M>1$, let
(77) $\mathcal{U}_{M}\coloneqq\left\{F\in C^{0}(X,\mathrm{GL}(d,\mathbb{F}))\;\mathord{:}\;\forall x\in X,\ \sup_{n\in\mathbb{Z}}\kappa(F^{(n)}(x))>M\right\}.$
A compactness argument shows that $\mathcal{U}_{M}$ is open in $C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$. We will establish that $\mathcal{U}_{M}$ is dense. From this it will follow that the set $\mathcal{QC}=\bigcap_{M}\mathcal{U}_{M}$ is meager, therefore completing the proof.
Instead of working with the metric (2), we use instead the metric
(78) $d^{\prime}(F,G)\coloneqq\sup_{x\in X}\|F(x)-G(x)\|\,,$
which, despite not being complete, induces the same topology on $C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$.
Fix $M>1$, $F\in C^{0}(X,\mathrm{GL}(d,\mathbb{F}))$, and $\varepsilon>0$. Let $C_{0}\coloneqq\sup_{x\in X}\|F(x)\|$. Fix $N>0$ such that
(79) $(1+\varepsilon)^{N}>M^{2}\,.$
Let $\delta>0$ be such that for all $x,y\in X$, if $d(x,y)<\delta$, then $\|F(x)-F(y)\|<\varepsilon$. By Theorem 4.1, there exists a castle partition
(80) $X=\bigsqcup_{i=1}^{s}\bigsqcup_{j=0}^{\ell_{i}-1}T^{j}B_{i}\,.$
where each tower $K_{i}\coloneqq\bigsqcup_{j=0}^{\ell_{i}}T^{j}B_{i}$ is $N$-capturing and each floor $T^{j}B_{i}$ is a clopen set of diameter $<\delta$.
For each $j\in\ldbrack 1,s\rdbrack$ and $j\in\ldbrack 0,\ell_{i}-1\rdbrack$, choose an arbitrary point $x_{i,j}$ in the set $T^{i}B_{j}$ and let $A_{i,j}\coloneqq F(x_{i,j})$. Using Lemmas 5.1 and 5.2, we find matrices $\tilde{A}_{i,j}$ such that $\|\tilde{A}_{i,j}-A_{i,j}\|\leq\varepsilon\|A_{i,j}\|$ and
(81) $\kappa_{\mathrm{e}}(\tilde{A}_{i,\ell_{i}-1}\cdots\tilde{A}_{i,0})$ $\geq(1+\varepsilon)^{\ell_{i}}$ if $\quad\text{if }\ell_{i}$ $<N\,,$
(82) $\kappa(\tilde{A}_{i,\ell_{i}-1}\cdots\tilde{A}_{i,0})$ $\geq(1+\varepsilon)^{\ell_{i}}$ if $\quad\text{if }\ell_{i}$ $\geq N\,.$
Now define a function $G\colon X\to\mathrm{GL}(d,\mathbb{F})$ as constant equal to $\tilde{A}_{i,j}$ on the set $T^{i}(B)$. Since those sets are clopen, $G$ is continuous.
Let us check that $G$ is a perturbation of $F$. Every point $x\in X$ belongs to a unique element $T^{j}(B_{i})$ of the castle partition. Then we can bound:
(83) $\|G(x)-F(x)\|$ $\leq\|\tilde{A}_{i,j}-A_{i,j}\|+\|F(x_{i,j})-F(x)\|$
(84) $<C\varepsilon+\varepsilon\,.$
Thus, $d^{\prime}(G,F)<(C+1)\varepsilon$.
The final task is checking that $G\in\mathcal{U}_{M}$, that is, for every $x\in M$,
(85) $\exists n\in\mathbb{Z}\text{ s.t.\ }\kappa(G^{(n)}(x))>M\,.$
So fix a point $x\in X$. Let $K_{i}$ be the tower containing $x$.
As a first case, assume that the height $\ell_{i}$ of the tower is at least $N$. We have $x=T^{j}y$ for some $y\in B_{i}$ and $j\in\ldbrack 0,\ell_{i}-1\rdbrack$. Then $\kappa(G^{(\ell_{i})}(y))\geq(1+\varepsilon)^{\ell_{i}}>M^{2}$. Since
(86) $G^{(\ell_{i})}(y)=G^{(\ell_{i}-j)}(x)\,[G^{(-j)}(x)]^{-1}\,,$
it follows from the sub-multiplicativity property (65) that either $\kappa(G^{(\ell_{i}-j)}(x))$ or $\kappa(G^{(-j)}(x))$ is bigger than $M$. This proves property (85) in this case.
Now assume that $\ell_{i}<N$. Since $K_{i}$ is an $N$-capturing set and also a tower of height $\ell_{i}$, it is actually a $m_{i}\ell_{i}$-capturing set, where $m_{i}\coloneqq\lceil N/\ell_{i}\rceil$. Using characterization (17) of the capturing property, we see that there exists a point $y\in B_{i}$ such that the segment of orbit $\{T^{n}y\;\mathord{:}\;n\in\ldbrack 0,m_{i}\ell_{i}-1\rdbrack\}$ is contained in $K_{i}$ and contains the point $x$. Then $\kappa(G^{(\ell_{i})}(y))\geq(1+\varepsilon)^{\ell_{i}}$; furthermore, $G^{(m_{i}\ell_{i})}(y)=[G^{(\ell_{i})}(y)]^{m_{i}}$, so
(87) $\kappa(G^{(m_{i}\ell_{i})}(y))\geq\kappa_{\mathrm{e}}(G^{(m_{i}\ell_{i})}(y))=[\kappa_{\mathrm{e}}(G^{(\ell_{i})}(y))]^{m_{i}}\geq(1+\varepsilon)^{m_{i}\ell_{i}}>M^{2}\,.$
Take $k\in\ldbrack 0,m_{i}\ell_{i}-1\rdbrack$ such that $T^{k}y=x$. Using sub-multiplicativity of $\kappa$ once again, we conclude that either $\kappa(G^{(m_{i}\ell_{i}-k)}(x))$ or $\kappa(G^{(-k)}(x))$ is bigger than $M$. This concludes the proof of (85). That is, $G\in\mathcal{U}_{M}$.
We have shown that the open set $\mathcal{U}_{M}$ is dense. The proof of Theorem 1.1 is complete. ∎
6. Further comments
The possibility of relaxing the hypotheses of Theorems 1.1 and 4.1 requires further investigation. It would be especially interesting to know if the results remain valid if we only require the space $X$ to have finite topological dimension. (Of course, the requirement that the floors are clopen in Theorem 4.1 would have to be relaxed as well.)
We note that Downarowicz’s [Do, Lemma 1] (which we mentioned in Section 2 above) admits an extension to spaces of any finite dimension, due to Gutman [G, Theorem 6.1]. However, Gutman’s theorem is false in infinite-dimensional spaces: see [TTY, Theorem 1.1]. Let us also point out a related tower theorem of Bonatti and Crovisier [BonC, Theorem 3.1], which applies to diffeomorphisms of finite-dimensional manifolds and allows for periodic points (with some technical restrictions).
Towers in finite-dimensional spaces have been used for perturbing linear cocycles in the papers [AvB, AvBD, Bo]. Some arguments from [AvBD] were later simplified without the use of towers: see [BoN]. Although we do not foresee a proof of Theorem 1.1 (or any substantive variation) that circumvents the use of towers, this may be a possibility worth considering.
Acknowledgements. I thank Meysam Nassiri, Hesam Rajabzadeh, and Zahra Reshadat for enlightening conversations. I am grateful to Sylvain Crovisier for bring reference [G] to my attention.